# Weak Lensing Approximation of Wave-optics Effects from General Symmetric Lens Profiles

### v1

@@ 6 Conclusion and Outlook
In this Letter, we have presented an analytical method to approximate the amplification factor $F(\omega,y)$ under the weak lensing conditions $\omega y^{2}/2\gg 1$ and $y\gg 1$ [^9] Orders of several tens are considered large in this work. The latter condition can be further relaxed if $y_{\mathrm{cr}}\lesssim 1$.. The approximation is particularly useful for moderately low $\omega$ with moderately large $y$, making it well-suited for studying WO effects caused by low-mass field halos and subhalos.
@@
Additionally, our method can be applied to composite lenses, such as a chameleon profile for baryonic matter embedded within a dark matter halo described by an NFW profile (Chiriv`ı et al., 2020), or an NFW lens with embedded subhalos (Zavala & Frenk, 2019). Furthermore, the method may be extended to non-symmetric lens profiles via perturbative techniques, provided the deviation of the lensing potential from axisymmetry is small—a condition thought to be common for galactic dark matter halos based on observational evidence (Meylan et al., 2006).
@@ Acknowledgments
O.A.H. acknowledges suport by grants from the Research Grants Council of Hong Kong (Project No. CUHK 14304622, 14307923, and 14307724), the start-up grant from the Chinese University of Hong Kong, and the Direct Grant for Research from the Research Committee of The Chinese University of Hong Kong. M. H. acknowledges support of the Science and Technology Facilities Council (Grant Ref. ST/V005634/1). Q. L. N. acknowledges support of the Institute for Cosmic Ray Research International University Research Program. This research has made use of NASA’s Astrophysics Data System Bibliographic Services.
@@
where $\mu$ is the magnification of the image. For symmetric lenses, $\Delta_{(1)}$ is given by Tambalo et al. (2023a); Savastano et al. (2023):

### v2

@@ 6 Conclusion and Outlook
In this paper, we have presented an analytical method to approximate the amplification factor $F(\omega,y)$ under the weak lensing conditions $\omega y^{2}/2\gg 1$ and $y\gg 1$ [^10] Orders of several tens are considered large in this work. The latter condition can be further relaxed if $y_{\mathrm{cr}}\lesssim 1$.. The approximation is particularly useful for moderately low $\omega$ with moderately large $y$, making it well-suited for studying WO effects caused by low-mass field halos and subhalos.
@@
Additionally, our method can be applied to nested symmetric lenses, such as a chameleon profile for baryonic matter embedded within a dark matter halo described by an NFW profile (Chirivì et al., 2020), or an NFW lens with embedded subhalos (Zavala & Frenk, 2019). Furthermore, the method may be extended to non-symmetric lens profiles via perturbative techniques, provided the deviation of the lensing potential from axisymmetry is small—a condition thought to be common for galactic dark matter halos based on observational evidence (Meylan et al., 2006).
@@ Acknowledgments
O.A.H. acknowledges support by grants from the Research Grants Council of Hong Kong (Project No. CUHK 14304622, 14307923, and 14307724), the start-up grant from the Chinese University of Hong Kong, and the Direct Grant for Research from the Research Committee of The Chinese University of Hong Kong. M. H. acknowledges support of the Science and Technology Facilities Council (Grant Ref. ST/V005634/1). Q. L. N. acknowledges support of the Institute for Cosmic Ray Research International University Research Program. This research has made use of NASA’s Astrophysics Data System Bibliographic Services.
@@
where $\mu$ is the magnification of the image. For symmetric lenses, $\Delta_{(1)}$ is given by (Tambalo et al., 2023a; Savastano et al., 2023):